OpenStudy (anonymous):

The y-intercept of the line whose equation is 2x + y = 4 is

OpenStudy (anonymous):

Do you know how to solve for y first?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

So, let's start by getting y by itself. If we are adding 2x to y, to get ride of it we need to subtract 2x from both sides of the equation like so: 2x + y = 4 -2x -2x _______________ What does that equal?

OpenStudy (anonymous):

y=4-2x?